# LLMs are Also Effective Embedding Models: An In-depth Overview

### v1

@@ 5 Tuning-based Embedding Model
While direct prompting of LLMs can yield useful embeddings, the true potential of these models is unlocked through a more refined approach: tuning them with existing or synthetic paired text data. This process typically employs contrastive learning to enhance the models’ ability to distinguish between semantically similar and dissimilar text pairs, resulting in more accurate and meaningful embeddings. For example, in the early stage of LLMs outbreak, Muennighoff (2022) initialized the embedding models with pre-trained GPT-3 models and applied continued contrastive training. The hidden state from the last layer corresponding to the special token [EOS] at the end of the sequence is taken as the embedding of the input sequence. Instructor (Su et al., 2023b) is the first work to explore a unified approach for generating text embeddings using task instructions based on GTR (Ni et al., 2021). It annotates instructions for 330 diverse tasks and trains the model on this multitask mixture with a contrastive loss.
@@
With the continuous progress of open-source LLMs, many researchers have attempted to use these models to build better text representation models. These advances primarily focus on optimizing model architecture (Lee et al., 2024b; Muennighoff et al., 2024), improving training objectives (BehnamGhader et al., 2024; Li et al., 2023c; Lee et al., 2024a, b), and further enhancing training data (Wang et al., 2024a; Kim et al., 2024; Jeronymo et al., 2023). In this section, we will delve into these three foundational techniques that underpin the fine-tuning of LLMs as powerful representation models. Table 2 summarize summarizes the configurations of representative models in terms of these three aspects.
5.1 Model Architecture
@@
The development of language models has seen significant changes in neural architecture design. Early models like BERT (Devlin et al., 2018) and ELMo (Peters et al., 2018) primarily utilized encoder-only architectures, which focus on bidirectional context, enabling the model to consider both left and right context during training. In contrast, recent LLMs have adopted decoder-only architectures (OpenAI, 2023; Touvron et al., 2023b; Jiang et al., 2023a), which are driven by the need for models to excel at generative tasks, such as text completion and creative writing. While encoder-only models provide strong contextual embeddings, decoder-only models have demonstrated superior performance in generative tasks due to their autoregressive nature.
@@
Embedding with Bi-directional Contextualization. Current generative LLMs predominantly use mono-directional attention, focusing on a unidirectional flow of information. This approach simplifies the model architecture and aligns well with autoregressive tasks where future tokens are predicted based on past context. However, the lack of bidirectional attention can limit the model’s ability to fully capture dependencies within the entire sequence. To address this, some models like Gecko (Lee et al., 2024b) and LLM2vec (BehnamGhader et al., 2024) propose incorporating bidirectional attention mechanisms within existing LLMs, enabling the model to consider both past and future tokens simultaneously. This enhancement aims to improve the quality of sequence embeddings by leveraging a more comprehensive understanding of the input text. In addition, GritLM (Muennighoff et al., 2024) unifies embedding tasks and generative tasks into a single model with bidirectional attention through generative representational instruction tuning. Recently, however, BGE-ICL (Li et al., 2024b) argues that enabling bidirectional attention during embedding fine-tuning misalign with the model’s original pre-training setup, which could compromise its effectiveness in generation tasks.
@@
Table 2: Detail of representative models that tune LLMs as embedding models. ‘MP’ is mean pooling operation. ‘sup.CL’ and ‘unsup.CL’ means supervised contrastive loss and unsupervised contrastive loss respectively. "Bi-Att" indicates whether the model enables bi-directional attention in LLMs, while "LoRA" specifies whether the model was trained using the LoRA technique.
Model Dim. # Params. Base Architecture Bi-Att LoRA Training Data Training Obj. (Neg)
SGPT (Muennighoff, 2022) 4,096 5.8B GPT Position-weighted MP $\times$ $\times$ SNLI/MNLI sup.CL
@@
GTE-Qwen2-7B-instruct (Li et al., 2023c) 3,584 7B Qwen2-7B MP $\checkmark$ $\times$ Unsup/Sup Pair unsup.CL & sup.CL
E5-mistral-7b-instruct (Wang et al., 2024a) 4,096 7B Mistral Last Token $\times$ $\checkmark$ E5(Public&Synthetic) sup.CL
@@
Echo-mistral-7b (Springer et al., 2024) 4,096 7B Mistral Last Token $\times$ $\checkmark$ E5 sup.CL
@@
SFR-Embedding 4,096 7B E5 Last Token $\times$ $\times$ Specially Curated Dataset sup.CL
NV-Embed (Lee et al., 2024a) 4,096 7B Mistral Latent Attention Layer $\checkmark$ $\checkmark$ Public Sup Datasets Two-stage sup.CL
Linq-Embed-Mistral (Kim et al., 2024) 4,096 7B E5 Last Token $\times$ $\checkmark$ E5S&Refined Synthetic sup.CL
Gecko (Lee et al., 2024b) 256 1.2B Transformer MP $\checkmark$ $\times$ FRet&Public Datasets sup.CL
@@
BGE-ICL (Li et al., 2024b) 4,096 7B Mistral Last Token $\times$ $\checkmark$ Public Sup Datasets sup.CL & KD
@@
Low-Rank Adaption. Traditional fine-tuning methods often require extensive computational resources and large amounts of labeled data. As a result, there is a growing interest in parameter-efficient tuning approaches (e.g., LoRA) that enable the adaptation of LLMs for embedding tasks while minimizing resource consumption (Wang et al., 2024a; Kim et al., 2024; Li et al., 2024b). These parameter-efficient tuning methods offer promising solutions by reducing the number of trainable parameters while maintaining comparable model performance or even better generalization performance (Wang et al., 2024a).
@@
Embedding Derivation. Obtaining sequence embeddings from these architectures involves different techniques. In BERT, the output states of [CLS] token is commonly used as a representation of the entire input sequence, providing a summary embedding. For LLMs, the [EOS] token, which signifies the end of a sequence, serves a similar purpose. Additionally, mean pooling of the last hidden layer can be employed to aggregate the contextual information across all tokens in the sequence (Su et al., 2023b; Springer et al., 2024; BehnamGhader et al., 2024; Lee et al., 2024b).Recent studies have explored more advanced methods to obtain sequence representations, such as using latent attention layer (Lee et al., 2024a) that adaptively combine token embeddings to enhance the final sequence representation, or sparse representations based on lexicon-importance distribution (Chen et al., 2024; Zhuang et al., 2024; Doshi et al., 2024; Lupart et al., 2024).
5.2 Training Objective
@@
Multi-task Training. Besides the standard contrastive objective, much effort has been made to improve the training procedures to enhance LLMs as versatile embedding models. For example, GTE (Li et al., 2023c) and Conan-embedding (Li et al., 2024e) introduce a multi-stage training approach in which the model is first pretrained with an InfoNCE loss using in-batch negatives on weakly supervised text relevance data, followed by fine-tuning with a CoSENT loss on the STS task. LLM2Vec (BehnamGhader et al., 2024) transforms a pre-trained decoder-only LLM into a universal text encoder through masked next token prediction and unsupervised contrastive learning. This method uses only publicly available data and applies unsupervised contrastive training similarly to SimCSE (Gao et al., 2021). NV-Embed (Lee et al., 2024a) also focuses on training procedures to enhance LLMs. Specifically, the model first performs contrastive training with instructions on retrieval datasets and then integrates carefully curated non-retrieval datasets into the stage-one training data. GritLM (Muennighoff et al., 2024) proposes a unified model for both embedding tasks and generative tasks, which is jointly optimized with both NLL objective and contrastive loss.
@@
Knowledge Distillation. Additionally, several studies explore using knowledge distillation to enhance embedding performance by leveraging larger embedding models or cross-attention ranker models. Gecko (Lee et al., 2024b), for instance, builds a smaller bidirectional embedding model (with 1.2B parameters) by distilling knowledge from a decoder-only LLM by generating synthetic paired data. BGE-ICL (Li et al., 2024b) introduce few-shot examples into the query side to enhance the query embeddings and also consider distilling the relevance score from the reranker for retrieval task during training. LLM Embedder introducs a unified embedding model designed to support the diverse retrieval augmentation needs of LLMs, including document knowledge, tools, in-context learning examples, and memory knowledge, through multi-task learning (Zhang et al., 2023b). The model incorporates a reward formulation based on LLM feedback and stabilizes knowledge distillation by integrating both soft reward-based labels and hard ranking-based labels during contrastive training. DiSCo proposes distilling the scores of rewritten queries paired with documents from a teacher model and utilizing knowledge distillation from multiple teachers to enhance conversational representations in conversational search tasks (Lupart et al., 2024).
@@
Table 3: Comparison of datasets utilized for various models. QuestionRR means StackOverFlowDupQuestionRR. -- indicates that the model uses this type of data, but specific datasets are not specified.
@@ 5.3 Data Construction
Training data plays a crucial role in the success of fine-tuning LLMs as embedding models. The methods typically employ two types of data: (1) adapting existing datasets, and (2) synthesizing datasets via LLMs.
@@
Adapting Existing Datasets. Table 3 presents the most commonly used public datasets, which include retrieval, clustering, classification, STS, and reranking tasks that align with the evaluation criteria in the Massive Text Embedding Benchmark (MTEB) [^1] https://huggingface.co/spaces/mteb/leaderboard (Muennighoff et al., 2022). These datasets provide diverse labeled examples that can be used to fine-tune models in a supervised manner, enhancing their ability to capture specific semantic relationships and improve performance on downstream embedding tasks. Specifically, these tasks are typically divided into symmetric and asymmetric tasks (Wang et al., 2024a). Symmetric tasks consist of queries and documents that convey similar meanings but differ in their wording, while asymmetric tasks include tasks where the query and document share a semantic relationship without being direct paraphrases. Besides paired data, some works leverage self-supervised learning methods like SimCSE, using natural language text from sources such as Wikipedia (BehnamGhader et al., 2024). These tasks often involve generating different views of the same input sequence through augmentations, such as dropout or token masking, to create positive pairs for contrastive learning. Moreover, SFR-Embedding-Mistral (Meng et al., 2024) fine-tunes the E5 model across a diverse array of tasks, and demonstrates that the effectiveness of embedding models can be enhanced through knowledge transfer from multiple tasks.
@@
Synthesizing Datasets via LLMs Considering the limited quantity of supervised data and the uneven distribution of corresponding pair relevance, recent research has addressed this challenge by leveraging LLMs to automatically synthesize large-scale, diverse, and contextually rich datasets (Wang et al., 2024a). This approach draws inspiration from LLM-based data augmentation in the supervised fine-tuning process (Xu et al., 2024). Synthetic datasets for training embedding models typically involve two primary methods: data labeling and data curation. In the labeling approach, a query is provided for the LLM to generate a relevant document as a positive example, optionally producing hard negative documents, or to generate a query for a sampled document from the corpus. In the curation method, the LLM is prompted to generate both the query and the relevant document, utilizing few-shot query-document examples as optional information. By using model-generated data, researchers can produce large-scale, high-quality training datasets tailored to specific embedding tasks, improving the robustness and generalizability of the fine-tuned models.
@@
In the early stage of LLM area, Inpars utilizes GPT-J to generate one synthetic query per given document, prompted with three examples from MS MARCO, and then leverage a powerful reranker to select the synthetic query-document pairs for training (jeronymo2023; Jeronymo et al., 2023). E5-Mistral introduces a two-step prompting method in which proprietary LLMs (GPT-4) are initially prompted to create a pool of candidate tasks (Wang et al., 2024a). From this pool, a specific task is chosen, and the LLMs are prompted again to generate triples consisting of a query, a positive document, and a hard-negative document. This method ultimately supports a broad range of text embedding tasks across 93 languages, covering hundreds of thousands of embedding tasks. Gecko (Lee et al., 2024b) prompts LLMs to read a sampled passage from a web corpus to generate both a task description and a relevant query, after which the top-K relevant passages were retrieved for these synthetic queries, and LLMs were employed to relabel more relevant positive passages along with a better hard negative. Linq-Embed-Mistral (Kim et al., 2024) utilizes GPT-4-turbo to generate the query-positive-negative triplet as training samples across six tasks, where the synthetic strategy was closely adhering to the (Wang et al., 2024a). Additionally, Linq conductS a data refinement on synthetic data to improve overall data quality through employing various types of few-shot prompt engineering, filtering, and negative mining (Kim et al., 2024). LARMOR explores the zero-shot capabilities of LLMs for generating queries, relevance judgments, and reference lists (Khramtsova et al., 2024).
@@
The selection method involves identifying hard negatives from the candidate corpus. Su et al. (2023b) employed Sentence-T5 embeddings (Ni et al., 2022) to encode pairs of texts, calculating pairwise cosine similarity scores to determine suitable hard negatives. Lee et al. (2024b) retrieved the top-K relevant passages for synthetic queries and utilized LLMs to relabel more relevant positive passages and identify better hard negatives. Li et al. (2024e) introduced dynamic hard negative mining, which selected more challenging negative examples during contrastive training. During the selection process, some studies have noted the issue of false negatives, where certain hard negatives should actually be classified as positive examples (Qu et al., 2020). To address this, Moreira et al. (2024) examined various filtering methods for hard negatives, proposing the use of similarity scores from positive examples as thresholds with specific margins or percentages.
@@
In contrast, the generation method leverages LLMs to directly create hard negatives for queries. Wang et al. (2024a) utilized LLMs to generate query, positive, hard-negative triples based on a selected task from pre-generted candidate tasks pool. This synthetic strategy has been applied to various embedding models (Muennighoff et al., 2024; Springer et al., 2024; Kim et al., 2024), demonstrating its versatility. Building on the approach of Wang et al. (2024a), Kim et al. (2024) introduced a data refinement process, such as such as few-shot prompt engineering, filtering, and improved negative mining, to enhance the quality of synthetic data. This integration of hard negative mining techniques directly exploits the powerful generative capabilities of LLMs, ultimately facilitates more robust representational learning, enhancing the capabilities of embedding models.
6 Advanced Embeddings

### v2

@@ 5. Tuning-based Embedding Models
While direct prompting of LLMs can yield useful embeddings, the true potential of these models is unlocked through a more refined approach: tuning them with existing or synthetic paired text data. This process typically employs contrastive learning to enhance the models’ ability to distinguish between semantically similar and dissimilar text pairs, resulting in more accurate and meaningful embeddings. For example, in the early stage of LLMs outbreak, (Muennighoff, 2022) initialized the embedding models with pre-trained GPT-3 models and applied continued contrastive training. The hidden state from the last layer corresponding to the special token [EOS] at the end of the sequence is taken as the embedding of the input sequence. Instructor (Su et al., 2023b) is the first work to explore a unified approach for generating text embeddings using task instructions based on GTR (Ni et al., 2021). It annotates instructions for 330 diverse tasks and trains the model on this multitask mixture with a contrastive loss.
@@
With the continuous progress of open-source LLMs, many researchers have attempted to use these models to build better text representation models. These advances primarily focus on optimizing model architecture (Lee et al., 2024a; Muennighoff et al., 2024), improving training objectives (BehnamGhader et al., 2024; Li et al., 2023d; Lee et al., 2024b, a), and further enhancing training data (Wang et al., 2024c; Kim et al., 2024; Jeronymo et al., 2023a). In this section, we will delve into these three foundational techniques that underpin the fine-tuning of LLMs as powerful representation models. Table 1 summarize summarizes the configurations of representative models in terms of these three aspects.
5.1. Model Architecture
Figure 3. An illustration of the evolution of text embeddings from ELMo and BERT to GPT-style models, from (Zaib et al., 2020).
@@ 5.1. Model Architecture
The development of language models has seen significant changes in neural architecture design. As illustrated in Figure 3, early models like ELMo (Peters et al., 2018) and BERT (Devlin et al., 2018) primarily utilized encoder-only architectures, which focus on bidirectional context, enabling the model to consider both left and right context during training. In contrast, recent LLMs have adopted decoder-only architectures (OpenAI, 2023; Touvron et al., 2023b; Jiang et al., 2023b), which are driven by the need for models to excel at generative tasks, such as text completion and creative writing. While encoder-only models provide strong contextual embeddings, decoder-only models have demonstrated superior performance in generative tasks due to their autoregressive nature.
@@
Embedding with Bi-directional Contextualization. Current generative LLMs predominantly use mono-directional attention, focusing on a unidirectional flow of information. This approach simplifies the model architecture and aligns well with autoregressive tasks where future tokens are predicted based on past context. However, the lack of bidirectional attention can limit the model’s ability to fully capture dependencies within the entire sequence. To address this, some models like Gecko (Lee et al., 2024a) and LLM2vec (BehnamGhader et al., 2024) propose incorporating bidirectional attention mechanisms within existing LLMs, enabling the model to consider both past and future tokens simultaneously. This enhancement aims to improve the quality of sequence embeddings by leveraging a more comprehensive understanding of the input text. In addition, GritLM (Muennighoff et al., 2024) unifies embedding tasks and generative tasks into a single model with bidirectional attention through generative representational instruction tuning. Recently, however, BGE-ICL (Li et al., 2024d) argue that enabling bidirectional attention during embedding fine-tuning misalign with the model’s original pre-training setup, which could compromise its effectiveness in generation tasks.
@@
Table 1. Detail of representative models that tune LLMs as embedding models. “sup.CL” and “unsup.CL” means supervised contrastive loss and unsupervised contrastive loss respectively. “Bi-Att” indicates whether the model enables bi-directional attention in LLMs, while “LoRA” specifies whether the model was trained using the LoRA technique. For the manner of embedding derivation, “MP” is mean pooling operation and “P-MP” means Position-weighted mean pooling, “LAT” means Latent Attention Layer.
Model Dim. # Params. Base Fusion Bi-Att LoRA Training Data Training Obj. (Neg)
SGPT (Muennighoff, 2022) 4,096 5.8B GPT P-MP $\times$ $\times$ SNLI/MNLI sup.CL
@@
GTE-Qwen2-7B (Li et al., 2023d) 3,584 7B Qwen2 MP $\checkmark$ $\times$ Unsup/Sup Pair unsup.CL & sup.CL
E5-mistral-7b (Wang et al., 2024c) 4,096 7B Mistral [EOS] $\times$ $\checkmark$ E5(Public&Synthetic) sup.CL
@@
Echo-mistral-7b (Springer et al., 2024) 4,096 7B Mistral [EOS] $\times$ $\checkmark$ E5 sup.CL
@@
SFR-Embedding 4,096 7B E5 [EOS] $\times$ $\times$ Specially Curated Dataset sup.CL
NV-Embed (Lee et al., 2024b) 4,096 7B Mistral LAT $\checkmark$ $\checkmark$ Public Sup Datasets Two-stage sup.CL
Linq-Embed-Mistral (Kim et al., 2024) 4,096 7B E5 [EOS] $\times$ $\checkmark$ E5S&Refined Synthetic sup.CL
Gecko (Lee et al., 2024a) 256 1.2B Transformer MP $\checkmark$ $\times$ FRet&Public Datasets sup.CL
@@
BGE-ICL (Li et al., 2024d) 4,096 7B Mistral [EOS] $\times$ $\checkmark$ Public Sup Datasets sup.CL & KD
Gemini-Embedding (Lee et al., [n. d.]) 3072 UNK Gemini MP $\times$ $\times$ Pub & Syn sup.CL
Qwen3-Embedding (Zhang et al., [n. d.]) 4096 4B/8B Qwen3 [EOS] $\times$ $\times$ Pub & Syn sup.CL
@@
Low-Rank Adaption. Traditional fine-tuning methods often require extensive computational resources and large amounts of labeled data. As a result, there is a growing interest in parameter-efficient tuning approaches (e.g., LoRA) that enable the adaptation of LLMs for embedding tasks while minimizing resource consumption (Wang et al., 2024c; Kim et al., 2024; Li et al., 2024d). These parameter-efficient tuning methods offer promising solutions by reducing the number of trainable parameters while maintaining comparable model performance or even better generalization performance (Wang et al., 2024c).
@@
Embedding Derivation. Obtaining sequence embeddings from these architectures involves different techniques. In BERT, the output states of [CLS] token is commonly used as a representation of the entire input sequence, providing a summary embedding. For LLMs, the [EOS] token, which signifies the end of a sequence, serves a similar purpose. Additionally, mean pooling of the last hidden layer can be employed to aggregate the contextual information across all tokens in the sequence (Su et al., 2023b; Springer et al., 2024; BehnamGhader et al., 2024; Lee et al., 2024a).Recent studies have explored more advanced methods to obtain sequence representations, such as using latent attention layer (Lee et al., 2024b) that adaptively combine token embeddings to enhance the final sequence representation, or sparse representations based on lexicon-importance distribution (Chen et al., 2024; Zhuang et al., 2024; Doshi et al., 2024; Lupart et al., 2024).
Figure 4. An overview of contextual expansion for embedding models. RMs means retrieval models.
Contextual Expansion. Compared with previous text embedding models which often only consider the text itself, some recent efforts expand the context of query or document to obtain enhanced text semantic information embedding. (Li et al., 2024d) incorporated few-shot examples on the query side to enhance query embeddings during supervised contrastive training. (Morris and Rush, [n. d.]) designed the alternative contrastive learning objective and the new contextual architecture to incorporate neighbor document information into document embeddings. (Yan et al., [n. d.]) generated thoughts towards the input query, then obtained the embedding of the query and generated thought separately, and finally incorporated both embeddings to produce the embedding that better reflects the query semantics. (Ji et al., [n. d.]) introduced the Chain-of-Deliberation mechanism, which encouraged LLM to encode documents with conducting reasoning process, enabling to generate more fine-grained document embeddings. (Neeser et al., [n. d.]) generated potential user queries based on each document, and through incorporating the embeddings of generated queries with the document embeddings, it’s more effective to capture diverse user queries that reference the same document in different ways. (Tejaswi et al., [n. d.]) proposed a simple yet effective method that enhances retriever models by training them with semantically similar in-context examples, enabling better utilization of contextual signals for improved retrieval performance. (Li et al., 2025b) proposed a Reinforced Information Retrieval framework in which the query-expansion model and embedding model mutually enhance each other. Expanded queries generate more informative embeddings, while refined embeddings guide more effective query expansions, creating a synergistic loop that improves retrieval accuracy. An overview of these contextual expansion techniques and their integration strategies within embedding-based retrieval models is illustrated in Figure 4.
Figure 5. An illustration of Matryoshka Embedding, from (Kusupati et al., 2022).
Matryoshka Embedding (ME). Inspired by Russian nesting dolls, where smaller parts nest within larger ones. Matryoshka Embeddings (Kusupati et al., 2022) are trained to store the most important information in the early dimensions of an embedding. This allows the embeddings to be truncated to smaller sizes (e.g., 64, 128, 256 dimensions) while still retaining useful information. As shown in Figure 5, by applying loss functions to multiple truncated versions during training, ME enables flexible trade-offs among performance, speed, and memory usage. It’s especially useful for tasks like retrieval or classification, where quick filtering can be done with smaller embeddings, followed by re-ranking with full-size ones. More recently, The Qwen3 Embedding model series (Zhang et al., [n. d.]) employs the Matryoshka Embedding training approach and offers flexible, user-defined output dimensions ranging from 32 to 4096. This enables the model to better balance performance and computational efficiency across different applications.
5.2. Training Objective
@@
Multi-task/stage Training. Besides the standard contrastive objective, much effort has been made to improve the training procedures to enhance LLMs as versatile embedding models. For example, GTE (Li et al., 2023d) and Conan-embedding (Li et al., 2024e) introduce a multi-stage training approach in which the model is first pretrained with an InfoNCE loss using in-batch negatives on weakly supervised text relevance data, followed by fine-tuning with a CoSENT loss on the STS task. LLM2Vec (BehnamGhader et al., 2024) transforms a pre-trained decoder-only LLM into a universal text encoder through masked next token prediction and unsupervised contrastive learning. This method uses only publicly available data and applies unsupervised contrastive training similarly to SimCSE (Gao et al., 2021). NV-Embed (Lee et al., 2024b) also focuses on training procedures to enhance LLMs. Specifically, the model first performs contrastive training with instructions on retrieval datasets and then integrates carefully curated non-retrieval datasets into the stage-one training data. GritLM (Muennighoff et al., 2024) proposes a unified model for both embedding tasks and generative tasks, which is jointly optimized with both NLL objective and contrastive loss. (Zhang et al., [n. d.]) also employs a multi-stage training pipeline, which includes weakly supervised training using large-scale synthetic paired data (over 150 million pairs), followed by supervised fine-tuning with high-quality synthetic and labeled data.
@@
Knowledge Distillation. Additionally, several studies explore using knowledge distillation to enhance embedding performance by leveraging larger embedding models or cross-attention ranker models. Gecko (Lee et al., 2024a), for instance, builds a smaller bidirectional embedding model (with 1.2B parameters) by distilling knowledge from a decoder-only LLM by generating synthetic paired data. BGE-ICL (Li et al., 2024d) introduce few-shot examples into the query side to enhance the query embeddings and also consider distilling the relevance score from the reranker for retrieval task during training. Zhang et al. (2023b) introduce a unified embedding model designed to support the diverse retrieval augmentation needs of LLMs, including document knowledge, tools, in-context learning examples, and memory knowledge, through multi-task learning. The model incorporates a reward formulation based on LLM feedback and stabilizes knowledge distillation by integrating both soft reward-based labels and hard ranking-based labels during contrastive training. Lupart et al. (2024) propose distilling the scores of rewritten queries paired with documents from a teacher model and utilizing knowledge distillation from multiple teachers to enhance conversational representations in conversational search tasks.
Figure 6. An illustration of various data construction methods for training LLM-based embedding models. Here, $d^{+}$ denotes the positive document, and $\{d^{-}\}$ represents the set of hard negative documents. The variable $m$ refers to diverse meta-information used to guide synthetic data generation. $I$ is the instruction provided to the LLM for generating either a query $q$ or a document $d$, while $c$ optionally includes few-shot demonstrations. Finally, $t$ denotes the task-specific instruction.
@@
Table 2. Comparison of datasets used for supervised fine-tuning of several representative embedding models. QuestionRR means StackOverFlowDupQuestionRR. “--” indicates that the model uses this type of data, but specific datasets are not specified.
@@ 5.3. Data Construction
Training data plays a crucial role in the success of fine-tuning LLMs as embedding models.The methods typically employ two types of data: (1) adapting existing datasets, and (2) synthesizing datasets via LLMs. The latter often involves both data curation and automated labeling to construct training pairs. Figure 6 illustrates representative methods for constructing contrastive training datasets under these paradigms.
@@
Adapting Existing Datasets. Table 2 presents the most commonly used public datasets, which include retrieval, clustering, classification, STS, and reranking tasks that align with the evaluation criteria in the Massive Text Embedding Benchmark (MTEB) [^1] https://huggingface.co/spaces/mteb/leaderboard (Muennighoff et al., 2022). These datasets provide diverse labeled examples that can be used to fine-tune models in a supervised manner, enhancing their ability to capture specific semantic relationships and improve performance on downstream embedding tasks. Specifically, these tasks are typically divided into symmetric and asymmetric tasks (Wang et al., 2024c). Symmetric tasks consist of queries and documents that convey similar meanings but differ in their wording, while asymmetric tasks include tasks where the query and document share a semantic relationship without being direct paraphrases. Besides paired data, some works leverage self-supervised learning methods like SimCSE, using natural language text from sources such as Wikipedia (BehnamGhader et al., 2024). These tasks often involve generating different views of the same input sequence through augmentations, such as dropout or token masking, to create positive pairs for contrastive learning. Moreover, SFR-Embedding-Mistral (Meng et al., 2024) fine-tunes the E5 model across a diverse array of tasks, and demonstrates that the effectiveness of embedding models can be enhanced through knowledge transfer from multiple tasks.
@@
Synthesizing Datasets via LLMs Considering the limited quantity of supervised data and the uneven distribution of corresponding pair relevance, recent research has addressed this challenge by leveraging LLMs to automatically synthesize large-scale, diverse, and contextually rich datasets (Wang et al., 2024c). This approach draws inspiration from LLM-based data augmentation in the supervised fine-tuning process (Xu et al., 2024). Synthetic datasets for training embedding models typically involve two primary methods: data labeling and data curation. In the labeling approach, a query is provided for the LLM to generate a relevant document as a positive example(Wang et al., 2024c), optionally producing hard negative documents(Wang et al., 2024c), or to generate a query(Zhang et al., [n. d.]; Lee et al., 2024a) for a sampled document from the corpus. In the curation method, the LLM is prompted to generate both the query and the relevant document, utilizing few-shot query-document examples as optional information. By using model-generated data, researchers can produce large-scale, high-quality training datasets tailored to specific embedding tasks, improving the robustness and generalizability of the fine-tuned models.
@@
In the early stage of LLM area, (Jeronymo et al., 2023b, a) utilize GPT-J to generate one synthetic query per given document, prompted with three examples from MS MARCO, and then leverage a powerful reranker to select the synthetic query-document pairs for training. E5-Mistral (Wang et al., 2024c) introduce a two-step prompting method in which proprietary LLMs (GPT-4) are initially prompted to create a pool of candidate tasks. From this pool, a specific task is chosen, and the LLMs are prompted again to generate triples consisting of a query, a positive document, and a hard-negative document. This method ultimately supports a broad range of text embedding tasks across 93 languages, covering hundreds of thousands of embedding tasks. Gecko (Lee et al., 2024a) prompted LLMs to read a sampled passage from a web corpus to generate both a task description and a relevant query, after which the top-K relevant passages were retrieved for these synthetic queries, and LLMs were employed to relabel more relevant positive passages along with a better hard negative. Linq-Embed-Mistral (Kim et al., 2024) utilized GPT-4-turbo to generate the query-positive-negative triplet as training samples across six tasks, where the synthetic strategy was closely adhering to the (Wang et al., 2024c). Additionally, (Kim et al., 2024) conducted a data refinement on synthetic data to improve overall data quality through employing various types of few-shot prompt engineering, filtering, and negative mining. (Khramtsova et al., 2024) explored the zero-shot capabilities of LLMs for generating queries, relevance judgments, and reference lists. (Lee et al., [n. d.]) leveraged multi-stage prompting strategies to generate training samples across retrieval and classification tasks. For retrieval, Gemini was prompted to generate queries for web passages followed by a Gemini auto-rater to filter lower-quality examples and for classification, counterfactual, sentiment, and review classification training samples were generated. (Zhang et al., [n. d.]) prompted Qwen3-32B to generate queries for the documents by employing diverse prompting strategies, where specific roles were assigned to each document with incorporating various generative dimensions including query type, length, difficulty, and language.
@@
The selection method involves identifying hard negatives from the candidate corpus. (Su et al., 2023b) employed Sentence-T5 embeddings (Ni et al., 2022) to encode pairs of texts, calculating pairwise cosine similarity scores to determine suitable hard negatives. (Lee et al., 2024a) retrieved the top-K relevant passages for synthetic queries and utilized LLMs to relabel more relevant positive passages and identify better hard negatives. (Li et al., 2024e) introduce dynamic hard negative mining, which selects more challenging negative examples during contrastive training. During the selection process, some studies have noted the issue of false negatives, where certain hard negatives should actually be classified as positive examples (Qu et al., 2020). To address this, (Moreira et al., 2024) examined various filtering methods for hard negatives, proposing the use of similarity scores from positive examples as thresholds with specific margins or percentages.
@@
In contrast, the generation method leverages LLMs to directly create hard negatives for queries. (Wang et al., 2024c) utilized LLMs to generate query, positive, hard-negative triples based on a selected task from pre-generted candidate tasks pool. This synthetic strategy has been applied to various embedding models (Muennighoff et al., 2024; Springer et al., 2024; Kim et al., 2024), demonstrating its versatility. Building on the approach of (Wang et al., 2024c), (Kim et al., 2024) introduced a data refinement process, such as such as few-shot prompt engineering, filtering, and improved negative mining, to enhance the quality of synthetic data. This integration of hard negative mining techniques directly exploits the powerful generative capabilities of LLMs, ultimately facilitates more robust representational learning, enhancing the capabilities of embedding models.
6. Specialized Embeddings